If we increase two opposite sides of a square by 4 feet and the other two opposite sides by 6 feet, we shall then have a rectangle whose area is 74 square feet more than the square. Find a side of the square.
Answer by mananth(16946)

side of square=x
area = x^2
length of rectangle = x+6
width of rectangle = (x+4)
Area = (x+4)(x+6)
(x+4)(x+6)-x^2=74
x^2+10x+24-x^2=74
10x=50
x=5